# Neutron star evolution by combining discontinuous Galerkin and finite volume methods

### v1

@@ III.2 Modified finite volume (FV) / finite difference (FD) method
For non-conservative systems we can then use the same discretization as in Eq. (28). However, we modify the differentiation matrix $D_{qr}^{\bar{i}}$ so that it corresponds to an FD method using only three points. At or close to the element boundary we use one-sided stencils, so that all derivatives can be computed locally, without information from neighboring elements. This results in a band diagonal matrix given by
@@
$-\frac{\sigma}{16h}\sum_{r=0}^{N}(d_{q_{1}r}^{\bar{1}}u_{rq_{2}q_{3}}+d_{q_{2}r}^{\bar{2}}u_{q_{1}rq_{3}}+d_{q_{3}r}^{\bar{3}}u_{q_{1}q_{2}r})$ (33)
to the RHS of Eq. (28). Here $\sigma$ is a dissipation factor that we usually choose to be of order $10^{-1}$. The matrix
@@
For conservative systems of the form (10) we use a FV scheme given by
[…]
@@
::=:= $\frac{\left[J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}(n^{\bar{1}}_{i}f^{i})^{*}\right]_{q_{1}+\frac{1}{2}}+\left[J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}(n^{\bar{1}}_{i}f^{i})^{*}\right]_{q_{1}-\frac{1}{2}}}{h},$
@@
Notice that the flux derivatives are written here in terms of values halfway between grid points, e.g. points $x^{\bar{1}}_{q_{1}+\frac{1}{2}\ q_{2}\ q_{3}}$ and $x^{\bar{1}}_{q_{1}-\frac{1}{2}\ q_{2}\ q_{3}}$. This means we have to obtain the flux $f^{i}$ at locations between grid points by interpolation. As one can see from Eq. (III.2), this interpolation can be done separately in each direction. Let us focus on the term in Eq. (LABEL:FV-term1). The numerical flux $(n^{\bar{1}}_{i}f^{i})^{*}$ in this expression is obtained by interpolating to the midpoints $x^{\bar{1}}_{q_{1}\pm\frac{1}{2}}$. For points far enough from the boundary we can use standard schemes such as WENO3 [98] or WENOZ [107]. Let us now consider $q_{1}=1$, i.e. one point in from the left boundary. In this case $(n^{\bar{1}}_{i}f^{i})^{*}_{1+\frac{1}{2}}$ can be obtained by simple WENO3 interpolation using data at the points with indices 0, 1, and 2. On the other hand, $(n^{\bar{1}}_{i}f^{i})^{*}_{1-\frac{1}{2}}$ is obtained by linear interpolation using data at the points with indices 0, and 1. This means that, near the boundary, our scheme is only second order accurate (error of $O(h^{2})$) because of the use of linear interpolation. Finally, we consider $q_{1}=0$, i.e. the point at the left boundary. In this case we first compute
@@
::=:= $\frac{\left[J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}(n^{\bar{1}}_{i}f^{i})^{*}\right]_{\frac{1}{2}}+\left[J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}(n^{\bar{1}}_{i}f^{i})^{*}\right]_{0}}{h/2}.$
Here $(n^{\bar{1}}_{i}f^{i})^{*}_{\frac{1}{2}}$ is again obtained by linear interpolation using data at the points with indices 0, and 1. For $(n^{\bar{1}}_{i}f^{i})^{*}_{0}$ the situation is different, as it is located directly at the boundary where we do have grid points. Here we use the same scheme as in the DG method and compute the numerical flux from the value in this element and the adjacent element.
@@
Figure 2: 1D Grid point arrangement for FV with WENO reconstruction. The numbering at the top shows the grid point $q_{1}$ = 0, 1, …. The dashed lines indicate the midpoints between grid points. The region between two consecutive midpoints is called a cell. For each cell we need to reconstruct fluxes at its left and right end. In the interior of the grid (in cells $q_{1}$ = 2, …, $N-2$), when we have 5 consecutive grid points available for interpolation, we use WENOZ reconstruction. In cells $q_{1}=1$ and $q_{1}=(N-1)$ we drop the order of interpolation and use WENO3 instead. One stencil for each of WENOZ and WENO3 has been shown in the figure to illustrate their actual usage of the grid points. Even closer to the boundary, e.g. in the cell between grid point 0 and midpoint $\frac{1}{2}$ (marked by $\times$), we perform linear interpolation using fields at $q_{1}=0$ and $q_{1}=1$ to obtain fluxes at midpoint $\frac{1}{2}$. The same is done at midpoint $N-\frac{1}{2}$. The fluxes at the endpoints ($q_{1}=0$ and $q_{1}=N$) can be directly computed from the fields there, without any interpolation. From the fluxes at the cell ends we construct a numerical flux as usual from data in this cell and the neighboring cell in the adjacent element. From the numerical fluxes at the cell ends we then obtain flux derivatives at the cell centers. For all interior cells the centers coincide with grid points, so that we immediately obtain flux derivatives at the grid points. However, the cell centers near the left and right end (marked by the $\times$) do not coincide with grid points. To obtain flux derivatives at grid points $q_{1}=0$ and $q_{1}=N$ we use a weighted average of the flux derivatives at the two closest cell centers of this element.
Figure 2 illustrates what type of interpolation we use to obtain values at the midpoints. Note however, that $\left[\frac{d(J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}n^{\bar{1}}_{i}f^{i})}{dx^{\bar{1}}}\right]_{\frac{1}{4}}$ corresponds to a derivative term at a point that is a quarter of a grid spacing away from the point at $q_{1}=0$. To obtain a value at $q_{1}=0$ we can simply copy this value to the point at $q_{1}=0$. However, this very simple procedure results in an error of $O(h)$ and may thus spoil second order accuracy. We therefore usually prefer to obtain the value at $q_{1}=0$ from
$\left[\frac{d(J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}n^{\bar{1}}_{i}f^{i})}{dx^{\bar{1}}}\right]_{0}$ ::=:= $w\left[\frac{d(J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}n^{\bar{1}}_{i}f^{i})}{dx^{\bar{1}}}\right]_{\frac{1}{4}}$
@@
i.e. from a weighted average of values at points $\frac{1}{4}$ and $1$. We usually choose $w=4/3$ to obtain linear extrapolation from points at $\frac{1}{4}$ and $1$ to 00. Note, however, that this kind of extrapolation has two drawbacks. First, it is not exactly conservative, but violations in mass conservation still converge to zero with increasing resolution. Second, extrapolation can become very inaccurate in low density regions (e.g. in atmosphere regions around the star). We therefore have introduced cutoffs below which we simply use $w=1$. For the cases where we activate these cutoffs, we set $w=1$ if either $\rho_{0}<10^{-9}\rho_{0,max}(t=0)$, or $\epsilon<10^{-9}\epsilon_{max}(t=0)$, or $\epsilon>10^{3}\epsilon_{max}(t=0)$, where $\rho_{0,max}(t=0)$ and $\epsilon_{max}(t=0)$ are the maximum values of $\rho_{0}$ and $\epsilon$ at the initial time.
@@ III.3 Time integration
Note that Eqs. (III.1), (III.1), and (III.2) still contain time derivatives, since up to this point we have only discretized spatial derivatives. This means that these equations represent a set of coupled ordinary differential equations for the fields $u_{q_{1}q_{2}q_{3}}$ at the grid points. In this paper, we use a standard strongly stability preserving third order Runge-Kutta (RK) time integration method [115] to find the solution of these ODEs. Such Runge-Kutta methods are only stable when the Courant-Friedrichs-Lewy (CFL) condition is satisfied, i.e., if the time step $\Delta t$ satisfies
@@ III.4 Filtering for improved stability
Even when the time step satisfies the CFL condition, instabilities can still occur in some cases in the DG elements, e.g., on non-Cartesian patches. To combat such instabilities, we filter out high frequency modes in the evolved fields in the DG elements. This is achieved by first computing the coefficients $c_{l_{1}l_{2}l_{3}}$ in the expansion
@@
III.5 Treatment of low density regions
In regions of low density, the hydrodynamic fields can develop unphysical values. This is due to numerical errors during RK substeps, and also due to the difficulties in recovering the primitive variables from the evolved conserved variables. To deal with these problems and to keep the fields in a physically sound state, we introduce two hydrodynamic numerical mechanisms. We introduce an artificial atmosphere in the vacuum and low density regions. Additionally, we apply some positivity limiters to the fields to further ensure physicality.
@@ III.5.1 Atmosphere treatment:
However, we treat the points in the intermediate density range of $\rho_{\mathrm{atmo,cut}}<\rho_{0}<\rho_{\mathrm{atmo,max}}$ with more caution, as these are the points in the low density range transitioning from the atmosphere to the bulk of the matter in the system. Such points are challenging for all numerical simulation methods. For these, simply setting $\rho_{0}=\rho_{\mathrm{atmo,level}}$, $\epsilon=0$ and $Wv^{i}=0$ will cause us to lose too much velocity and energy from the system, and deviate from the correct GRHD solution. At the same time, we have to prevent the particles in the transitional density region from gaining arbitrarily large velocity and energy due to accumulating numerical errors. For this purpose, we appropriately choose a maximum cutoff for the magnitude of $Wv=\displaystyle\sqrt{Wv_{i}Wv^{i}}$ in the atmosphere regions as $(Wv)_{\mathrm{atmo,max}}=0.01$, and another cutoff for the $\epsilon$ magnitude in the atmosphere as $\epsilon_{\mathrm{atmo,max}}=0.1$, Then, in this transition density range $\rho_{\mathrm{atmo,cut}}<\rho_{0}<\rho_{\mathrm{atmo,max}}$, if at any point we find $(Wv)^{2}>(Wv)^{2}_{\mathrm{atmo,max}}$ we scale the $Wv^{i}$ components at that point as $Wv^{i}\rightarrow\displaystyle Wv^{i}\cdot\sqrt{\frac{(Wv)^{2}_{\mathrm{atmo,max}}}{(Wv)^{2}}}$. Additionally, if $\epsilon$ goes above $\epsilon_{\mathrm{atmo,max}}$, we simply reset $\epsilon$ to $\epsilon_{\mathrm{atmo,max}}$. The density $\rho_{0}$ is not modified in the transitional region.
@@ III.5.2 Positivity limiters:
To prevent our hydrodynamical fields from getting unphysical, we use a number of positivity limiters. These mainly target low density regions, but can activate in any region where unphysicality occur. In this context, we consider “unphysical” if (i) mass density $D$ becomes negative, (ii) energy density $\tau$ becomes negative, or (iii) if $S>D+\tau$, where $S=\sqrt{S_{i}S^{i}}$.
@@
where $S=\displaystyle\sqrt{S_{i}S^{i}}$ and the $10^{-14}$ subtraction ensures that $S$ remains below $D+\tau$. The $\min$ in Eq. (47) refers to the minimum over all points in the element. We use $\theta_{S}$ for scaling all the $S_{i}$ values in the element. For $D$ and $\tau$ related limiting, if the denominator becomes 00, we simply set $\theta_{u}=1$, and for $S_{i}$ we set it to $1-10^{-14}$. Lastly, $\theta_{u}$ is always constrained within the range $0\leq\theta_{u}\leq 1$. Note that all these limiting strategies preserve the element averages of the variables, which ensure better accuracy for DG elements. However, if the average itself violates the condition we try to enforce, the above scaling towards the average is not sufficient to obtain a physical state. In order to address this eventuality, we also perform a pointwise limiting for all elements to fix the state of the fields in any problematic points. The checked conditions still remain the same as (i) through (iii) as stated above. In the pointwise limiting, in case at a point $S>D+\tau$, we simply scale $S_{i}$ as $S_{i}\to\displaystyle S_{i}\cdot\left|\frac{D+\tau}{S}\right|\cdot f_{S}^{\mathrm{lim}}$, where we choose $f_{S}^{\mathrm{lim}}$ appropriately to be $0.99$. In case we find $D<0$ or $D+\tau\leq 0$, we simply set $S_{i}=0$. As for $D$, if at any point we find $D<10^{-12}\rho_{0,\mathrm{max}}(t=0)$, we simply set $D=10^{-12}\rho_{0,\mathrm{max}}(t=0)$. Similarly, for $\tau$, if at any point we find $\tau<0$, we set $\tau=0$.
@@
In addition to limiting the variables $(D,\tau,S_{i})$ in an attempt to retain physicality, we also collect information regarding the field conditions in the non-atmosphere regions during each of our time evolution substeps. This is because if the unphysicallities arise at any point in an element and limiting becomes necessary, it is treated as sign of “trouble” in an element as the matter fields are not well-behaved in it. This information is then used in our list of criteria for determining trouble in an element discussed in Sec. III.7.

### v2

@@ III.2 Modified finite volume / finite difference method
For nonconservative systems we can then use the same discretization as in Eq. (III.1). However, we modify the differentiation matrix $D_{qr}^{\bar{i}}$ so that it corresponds to an FD method using only three points. At or close to the element boundary we use one-sided stencils, so that all derivatives can be computed locally, without information from neighboring elements. This results in a band diagonal matrix  [^1] On its own this matrix does not possess the summation by parts property discussed in e.g. [114]. However, Eq. (III.1) has extra terms on its right-hand side, which act as penalty terms, that help to stabilize the method. given by
@@
$\frac{\sigma}{16h}\sum_{r=0}^{N}(d_{q_{1}r}^{\bar{1}}u_{rq_{2}q_{3}}+d_{q_{2}r}^{\bar{2}}u_{q_{1}rq_{3}}+d_{q_{3}r}^{\bar{3}}u_{q_{1}q_{2}r})$ (33)
to the right-hand side (RHS) of Eq. (III.1). Here $\sigma$ is a dissipation factor that we usually choose to be of order $10^{-1}$. The matrix
@@
For conservative systems of the form (10) we use an FV scheme given by
$\partial_{t}u_{q_{1}q_{2}q_{3}}+\frac{1}{J_{q_{1}q_{2}q_{3}}}\Bigg\{$ $\left[\frac{d(J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}n^{\bar{1}}_{i}f^{i})}{dx^{\bar{1}}}\right]_{q_{1}q_{2}q_{3}}$
@@
$:=$ $\frac{\left[J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}(n^{\bar{1}}_{i}f^{i})^{*}\right]_{q_{1}+\frac{1}{2}}+\left[J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}(n^{\bar{1}}_{i}f^{i})^{*}\right]_{q_{1}-\frac{1}{2}}}{h},$
@@
Notice that the flux derivatives are written here in terms of values halfway between grid points, e.g. points $x^{\bar{1}}_{q_{1}+\frac{1}{2}\ q_{2}\ q_{3}}$ and $x^{\bar{1}}_{q_{1}-\frac{1}{2}\ q_{2}\ q_{3}}$. This means we have to obtain the flux $f^{i}$ at locations between grid points by interpolation. As one can see from Eq. (III.2), this interpolation can be done separately in each direction. Let us focus on the term in Eq. (LABEL:FV-term1). The numerical flux $(n^{\bar{1}}_{i}f^{i})^{*}$ in this expression is obtained by interpolating to the midpoints $x^{\bar{1}}_{q_{1}\pm\frac{1}{2}}$. For points far enough from the boundary we can use standard schemes such as WENO3 [98] or WENOZ [107]. Let us now consider $q_{1}=1$, i.e. one point in from the left boundary. In this case $(n^{\bar{1}}_{i}f^{i})^{*}_{1+\frac{1}{2}}$ can be obtained by simple WENO3 interpolation using data at the points with indices 0, 1, and 2. On the other hand, $(n^{\bar{1}}_{i}f^{i})^{*}_{1-\frac{1}{2}}$ is obtained by linear interpolation using data at the points with indices 0 and 1. This means that, near the boundary, our scheme is only second order accurate (error of $O(h^{2})$) because of the use of linear interpolation. Finally, we consider $q_{1}=0$, i.e. the point at the left boundary. In this case we first compute
@@
$:=$ $\frac{\left[J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}(n^{\bar{1}}_{i}f^{i})^{*}\right]_{\frac{1}{2}}+\left[J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}(n^{\bar{1}}_{i}f^{i})^{*}\right]_{0}}{h/2}.$
Here $(n^{\bar{1}}_{i}f^{i})^{*}_{\frac{1}{2}}$ is again obtained by linear interpolation using data at the points with indices 0 and 1. For $(n^{\bar{1}}_{i}f^{i})^{*}_{0}$ the situation is different, as it is located directly at the boundary where we do have grid points. Here we use the same scheme as in the DG method and compute the numerical flux from the value in this element and the adjacent element.
@@
Figure 2: 1D grid point arrangement for FV with WENO reconstruction. The numbering at the top shows the grid point $q_{1}$ = 0, 1, …. The dashed lines indicate the midpoints between grid points. The region between two consecutive midpoints is called a cell. For each cell we need to reconstruct fluxes at its left and right end. In the interior of the grid (in cells $q_{1}$ = 2, …, N2N-2), when we have five consecutive grid points available for interpolation, we use WENOZ reconstruction. In cells $q_{1}=1$ and $q_{1}=(N-1)$ we drop the order of interpolation and use WENO3 instead. One stencil for each of WENOZ and WENO3 has been shown in the figure to illustrate their actual usage of the grid points. Even closer to the boundary, e.g. in the cell between grid point 0 and midpoint $\frac{1}{2}$ (marked by $\times$), we perform linear interpolation using fields at $q_{1}=0$ and $q_{1}=1$ to obtain fluxes at midpoint $\frac{1}{2}$. The same is done at midpoint $N-\frac{1}{2}$. The fluxes at the endpoints ($q_{1}=0$ and $q_{1}=N$) can be directly computed from the fields there, without any interpolation. From the fluxes at the cell ends we construct a numerical flux as usual from data in this cell and the neighboring cell in the adjacent element. From the numerical fluxes at the cell ends we then obtain flux derivatives at the cell centers. For all interior cells the centers coincide with grid points, so that we immediately obtain flux derivatives at the grid points. However, the cell centers near the left and right end (marked by the $\times$) do not coincide with grid points. To obtain flux derivatives at grid points $q_{1}=0$ and $q_{1}=N$ we use a weighted average of the flux derivatives at the two closest cell centers of this element.
Figure 2 illustrates what type of interpolation we use to obtain values at the midpoints. Note however, that $\left[\frac{d(J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}n^{\bar{1}}_{i}f^{i})}{dx^{\bar{1}}}\right]_{\frac{1}{4}}$ corresponds to a derivative term at a point that is a quarter of a grid spacing away from the point at $q_{1}=0$. To obtain a value at $q_{1}=0$ we can simply copy this value to the point at $q_{1}=0$. However, this very simple procedure results in an error of $O(h)$ and may thus spoil second order accuracy. We therefore prefer to obtain the value at $q_{1}=0$ from
$\left[\frac{d(J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}n^{\bar{1}}_{i}f^{i})}{dx^{\bar{1}}}\right]_{0}$ $:=$ $w\left[\frac{d(J\sqrt{\bar{\gamma}^{\bar{1}\bar{1}}}n^{\bar{1}}_{i}f^{i})}{dx^{\bar{1}}}\right]_{\frac{1}{4}}$
@@
i.e., from a weighted average of values at points $\frac{1}{4}$ and $1$. We usually choose $w=4/3$ to obtain linear extrapolation from points at $\frac{1}{4}$ and $1$ to 0. Note, however, that this kind of extrapolation has two drawbacks. First, it is not exactly conservative, but violations in mass conservation still converge to zero with increasing resolution. Second, extrapolation can become very inaccurate in low-density regions (e.g. in atmosphere regions around the star). We therefore have introduced cutoffs below which we simply use $w=1$. For the cases where we activate these cutoffs, we set $w=1$ if either $\rho_{0}<10^{-9}\rho_{0,max}(t=0)$, or $\epsilon<10^{-9}\epsilon_{max}(t=0)$, or $\epsilon>10^{3}\epsilon_{max}(t=0)$, where $\rho_{0,max}(t=0)$ and $\epsilon_{max}(t=0)$ are the maximum values of $\rho_{0}$ and $\epsilon$ at the initial time.
@@ III.3 Time integration
Note that Eqs. (III.1), (III.1), and (III.2) still contain time derivatives, since up to this point we have only discretized spatial derivatives. This means that these equations represent a set of coupled ordinary differential equations for the fields $u_{q_{1}q_{2}q_{3}}$ at the grid points. In this paper, we use a standard strongly stability-preserving third-order Runge-Kutta (RK) time integration method [116] to find the solution of these ordinary differential equations (ODEs). Such Runge-Kutta methods are only stable when the Courant-Friedrichs-Lewy (CFL) condition is satisfied, i.e., if the time step $\Delta t$ satisfies
@@ III.4 Filtering for improved stability
Even when the time step satisfies the CFL condition, aliasing-driven instabilities can still occur in some cases in the DG elements, e.g., on non-Cartesian patches. To combat such instabilities, we filter out high-frequency modes in the evolved fields in the DG elements. This is achieved by first computing the coefficients $c_{l_{1}l_{2}l_{3}}$ in the expansion
@@
III.5 Treatment of low-density regions
In regions of low density, the hydrodynamic fields can develop unphysical values. This is due to numerical errors during RK substeps, and also due to the difficulties in recovering the primitive variables from the evolved conserved variables. To deal with these problems and to keep the fields in a physically sound state, we introduce two hydrodynamic numerical mechanisms. We introduce an artificial atmosphere in the vacuum and low-density regions. Additionally, we apply some positivity limiters to the fields to further ensure physicality.
@@ III.5.1 Atmosphere treatment:
However, we treat the points in the intermediate density range of $\rho_{\mathrm{atmo,cut}}<\rho_{0}<\rho_{\mathrm{atmo,max}}$ with more caution, as these are the points in the low-density range transitioning from the atmosphere to the bulk of the matter in the system. Such points are challenging for all numerical simulation methods. For these, simply setting $\rho_{0}=\rho_{\mathrm{atmo,level}}$, $\epsilon=0$ and $Wv^{i}=0$ will cause us to lose too much velocity and energy from the system, and deviate from the correct GRHD solution. At the same time, we have to prevent the particles in the transitional density region from gaining arbitrarily large velocity and energy due to accumulating numerical errors. For this purpose, we appropriately choose a maximum cutoff for the magnitude of $Wv=\displaystyle\sqrt{Wv_{i}Wv^{i}}$ in the atmosphere regions as $(Wv)_{\mathrm{atmo,max}}=0.01$, and another cutoff for the $\epsilon$ magnitude in the atmosphere as $\epsilon_{\mathrm{atmo,max}}=0.1$. Then, in this transition density range $\rho_{\mathrm{atmo,cut}}<\rho_{0}<\rho_{\mathrm{atmo,max}}$, if at any point we find $(Wv)^{2}>(Wv)^{2}_{\mathrm{atmo,max}}$ we scale the $Wv^{i}$ components at that point as $Wv^{i}\rightarrow\displaystyle Wv^{i}\cdot\sqrt{\frac{(Wv)^{2}_{\mathrm{atmo,max}}}{(Wv)^{2}}}$. Additionally, if $\epsilon$ goes above $\epsilon_{\mathrm{atmo,max}}$, we simply reset $\epsilon$ to $\epsilon_{\mathrm{atmo,max}}$. The density $\rho_{0}$ is not modified in the transitional region.
@@ III.5.2 Positivity limiters:
To prevent our hydrodynamical fields from getting unphysical, we use a number of positivity limiters. These mainly target low-density regions, but can activate in any region where unphysicality occurs. In this context, we consider “unphysical” if (i) mass density $D$ becomes negative, (ii) energy density $\tau$ becomes negative, or (iii) if $S>D+\tau$, where $S=\sqrt{S_{i}S^{i}}$.
@@
where $S=\displaystyle\sqrt{S_{i}S^{i}}$ and the $10^{-14}$ subtraction ensures that $S$ remains below $D+\tau$. The $\min$ in Eq. (47) refers to the minimum over all points in the element. We use $\theta_{S}$ for scaling all the $S_{i}$ values in the element. For $D$ and $\tau$ related limiting, if the denominator becomes 0, we simply set $\theta_{u}=1$, and for $S_{i}$ we set it to $1-10^{-14}$. Lastly, $\theta_{u}$ is always constrained within the range $0\leq\theta_{u}\leq 1$. Note that all these limiting strategies preserve the element averages of the variables, which ensure better accuracy for DG elements. However, if the average itself violates the condition we try to enforce, the above scaling towards the average is not sufficient to obtain a physical state. In order to address this eventuality, we also perform a pointwise limiting for all elements to fix the state of the fields in any problematic point. The checked conditions still remain the same as (i) through (iii) as stated above. In the pointwise limiting, in case at a point $S>D+\tau$, we simply scale $S_{i}$ as $S_{i}\to\displaystyle S_{i}\cdot\left|\frac{D+\tau}{S}\right|\cdot f_{S}^{\mathrm{lim}}$, where we choose $f_{S}^{\mathrm{lim}}$ appropriately to be $0.99$. In case we find $D<0$ or $D+\tau\leq 0$, we simply set $S_{i}=0$. As for $D$, if at any point we find $D<10^{-12}\rho_{0,\mathrm{max}}(t=0)$, we simply set $D=10^{-12}\rho_{0,\mathrm{max}}(t=0)$. Similarly, for $\tau$, if at any point we find $\tau<0$, we set $\tau=0$.
@@
In addition to limiting the variables $(D,\tau,S_{i})$ in an attempt to retain physicality, we also collect information regarding the field conditions in the nonatmosphere regions during each of our time evolution substeps. This is because if the unphysicalities arise at any point in an element and limiting becomes necessary, it is treated as sign of “trouble” in an element as the matter fields are not well behaved in it. This information is then used in our list of criteria for determining trouble in an element discussed in Sec. III.7.